OpenStudy (anonymous):

solve the equation. 2sinx-1=0

OpenStudy (anonymous):

Set the equations so that sinx is by itself. Then solve for x. 2sinx - 1 = 0 2sinx = 1 sinx = 1/2 x = ???

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

http://www.mathsisfun.com/geometry/images/circle-unit-304560.gif Find the degree in which sinx = 1/2 sine is the y coordinate.